a nurse mix 15 liters of a 10% solution of s drug with som 60% solution to get a 20% MIXTURE. how many liters of the 60% solution will be needed?
Answer by checkley77(12844)

.60x+.10*15=.20(15+x)
.60x+1.5=3+.20x
.60x-.20x=3-1.5
.40x=1.5
x=1.5/.4
x=3.75 liters of 60% solution is used.
Proof:
.60*3.75+.10*15=.20(15+3.75)
2.25+1.5=.20*18.75
3.75=3.75